OpenStudy (anonymous):

A woman stands on a scale in a moving elevator. Her mass is 59.8 kg, and the combined mass of the elevator and scale is an additional 830 kg. Starting from rest, the elevator accelerates upward. During the acceleration, the hoisting cable applies a force of 9300 N. What does the scale read during the acceleration? A woman stands on a scale in a moving elevator. Her mass is 59.8 kg, and the combined mass of the elevator and scale is an additional 830 kg. Starting from rest, the elevator accelerates upward. During the acceleration, the hoisting cable applies a force of 9300 N. What does the scale read during the acceleration? @Mathematics

OpenStudy (anonymous):

The acceleration of the system is a = m/F = 889.8/9300 = 0.096 m/s^2 Now we now that the normal force is equal to the reading of the scale and our free body diagram shows that ma = Fn - Fg 59.8*0.096 = Fn - 59.8*9.8 Fn = 591.76 N or the mass according to the scale is Fn/g = 90.4 kg

OpenStudy (anonymous):

computer is not accepting the unit kg

OpenStudy (anonymous):

saying no conversion found between kg and the required unit

OpenStudy (anonymous):

The scale might be reading the normal force so use that

OpenStudy (anonymous):

WHICH WILL BE?

OpenStudy (anonymous):

591.76N

OpenStudy (anonymous):

IS INCORRECT

OpenStudy (anonymous):

whoops I forgot to take gravity into account for the acceleration of the system. This yields an acceleration of 1.53 m/^2 and a normal force of 677.79

OpenStudy (anonymous):

so the answe is what now?